OpenStudy (konradzuse):

find the derivative of y = arctan(x/9) − 1/(2(x2 + 81) )

OpenStudy (konradzuse):

d/dx of arctan(x/9) = \[\frac{\frac{1}{9}}{1+(\frac{x}{9})^2}\]

OpenStudy (konradzuse):

\[\frac{x}{(x^2+81)^2}\] for the other one...

OpenStudy (turingtest):

looks good

OpenStudy (konradzuse):

that can't be the answer tho..

OpenStudy (turingtest):

why not?

OpenStudy (konradzuse):

Idk maybe it cna, I was self checking with wolfman and it's dioofferent.... http://www.wolframalpha.com/input/?i=find+the+derivative+of+y+%3D+arctan%28x%2F9%29++%E2%88%92++1%2F%282%28x2+%2B+81%29++%29

OpenStudy (turingtest):

it's just some algebra; a different way of writing the same thing

OpenStudy (turingtest):

try simplifying the first expression

OpenStudy (konradzuse):

it could be 1+(x^2/81?)

OpenStudy (konradzuse):

1/9/that :P

OpenStudy (turingtest):

\[\frac{\frac{1}{9}}{1+(\frac{x}{9})^2}\cdot\frac99=\frac1{9+9x^2}\]so we have\[\frac1{9+9x^2}+\frac x{(x^2+81)^2}\]then we gotta rationalize that

OpenStudy (agent47):

Happy 100th anniversary Turing!

OpenStudy (turingtest):

thanks, lol

OpenStudy (konradzuse):

turning is 100? no wonder :P.

OpenStudy (turingtest):

ooh I should have multiplied by 9^2=81 I think\[\frac{81}{81+{x^2}}+\frac x{(x^2+81)^2}={(x^2+81)^2+x(81+x^2)\over(81+x^2)^2}\]

OpenStudy (turingtest):

with me so far?

OpenStudy (konradzuse):

mhmhm

OpenStudy (turingtest):

we want the denominator on the first to look like that of the other, and we can do that by multiplying by \(\frac{9^2}{9^2}=\frac{81}{81}\)\[{\frac19\over1+(\frac x9)^2}\cdot\frac{81}{81}={9\over81+x^2}\]oh, I made a mistake earlier... so that is the first term

OpenStudy (konradzuse):

extra hairy....

OpenStudy (konradzuse):

I think I got it from the wolfman explanation... I don't want you to waste your time :). Thanks!

OpenStudy (turingtest):

so now we have\[{9\over81+x^2}+{x\over(x^2+81)^2}\]

OpenStudy (turingtest):

welcome!

OpenStudy (konradzuse):

onto the next questions hehehe :).